A pair of dice is loaded. The probability that a 4 appears on the first die is 2/7, and the


probability that a 3 appears on the second die is 2/7. Other outcomes for each die appear


with probability 1/7. What is the probability of 7 appearing as the sum of the numbers


when the two dice are rolled?

1
Expert's answer
2022-02-23T10:04:00-0500

First die


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 1 & 2 & 3 & 4 & 5 & 6\\ \hline
 p(x) & 1/7 & 1/7 & 1/7 & 2/7 & 1/7 & 1/7\\

\end{array}"

Second die


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 1 & 2 & 3 & 4 & 5 & 6\\ \hline
 p(x) & 1/7 & 1/7 & 2/7 & 1/7 & 1/7 & 1/7\\

\end{array}"

"1+6=2+5=3+4=4+3=5+2=6+1"

"P(sum=7)=\dfrac{1}{7}(\dfrac{1}{7})+\dfrac{1}{7}(\dfrac{1}{7})+\dfrac{1}{7}(\dfrac{1}{7})"

"+\dfrac{2}{7}(\dfrac{2}{7})+\dfrac{1}{7}(\dfrac{1}{7})+\dfrac{1}{7}(\dfrac{1}{7})=\dfrac{9}{49}"